OpenStudy (anonymous):

A survey indicated that 3 out of 5 doctors use brand X aspirin. If 2000 doctors were surveyed, how many used brand X?

OpenStudy (anonymous):

1200 doctors used brand X

OpenStudy (anonymous):

3/5 = 2000 3*2000 = 6000 6000/5 = 1200 1200/2000 = 0.6 3/5 = 0.6

OpenStudy (anonymous):

@sgilmore208 Hope that helps